If 5 melons sell for $0.65, how much would 6 melons cost?
Answer by stanbon(75887)

If 5 melons sell for $0.65, how much would 6 melons cost?
----
Use a proportion:
x/6 = 0.65/5
---
x = (6/5)0.65
---
x = 6*0.13
---
x = 0.78 (cost of 6 melons)
Cheers,
Stan H.